What is the elapsed time for a 3rd observer in the Twin Paradox who remains stationary with respect to the centre of mass? My conjecture is t3 = gamma1 * t2 = gamma2 * t2, with velocities reckoned with respect to the 3rd observer.
1
Expert's answer
2019-04-29T09:23:03-0400

Assume that the moving observer A has mass "m_A" and moves with speed "v". Then let the stationary observer S which is on the Earth have mass "m_S" at 0 speed and let the beginning of x-axis be at the Earth as well. The position of center of mass is thus


"x_C=\frac{x_Am_A+0\cdot m_S}{m_A+m_S},"

where "x_A" - position of the moving observer A. This point is a function of time:


"x_A=v_A\cdot t,"

where "t" - time for the stationary observer S. Remember now that our observer A moves with very high speed, thus his mass will increase and finally the position of center of mass will be the following distance from earth:


"x_C=\frac{v_At\frac{m_A}{\sqrt{1-v^2_A/c^2}}}{\frac{m_A}{\sqrt{1-v^2_A/c^2}}+m_S},"


to simplify this expression, use Lorentz factor:


"\gamma_A=\frac{1}{\sqrt{1-v^2_A/c^2}},"

which gives


"x_C=\frac{v_At\cdot m_A\gamma_A}{m_A\gamma_A+m_S}."

Now put our middle observer C at the center of mass. As we see, the center of mass is a function of time. The more time the experiment lasts - the further from earth not only A, but C also becomes. What is the speed of C?


"v_C=\frac{x_C}{t}=v_A\frac{m_A\gamma_A}{m_A\gamma_A+m_S}."

So for this observer the elapsed time will be


"t_C=t_3=t\sqrt{1-\frac{v_C^2}{c^2}}=t\sqrt{1-\frac{v_A^2}{c^2}\Big(\frac{m_A\gamma_A}{m_A\gamma_A+m_S}\Big)^2}."

In case "m_S=m_A":


"t_C=t_3=t\sqrt{1-\frac{v_A^2}{c^2}\frac{\gamma_A^2}{(\gamma_A+1)^2}}."

Now express the ratio "v_A^2\/c^2" in terms of "\gamma_A" and substitute this to time C equation above:


"t_C=t\sqrt{1-\frac{\gamma_A^2-1}{\gamma_A^2}\frac{\gamma_A^2}{(\gamma_A+1)^2}}=t\sqrt{1-\frac{\gamma_A^2-1}{(\gamma_A+1)^2}}=\sqrt{\frac{2}{\gamma_A+1}}."